OpenStudy (anonymous):

The Digit in the unit place of the product 127^79 * 2^24?

OpenStudy (anonymous):

@UnkleRhaukus

OpenStudy (anonymous):

plzz help

OpenStudy (unklerhaukus):

can i use a calculator?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

any tricks???

OpenStudy (unklerhaukus):

calculator didn't help at all

OpenStudy (anonymous):

i think there is some tricks to do it....

OpenStudy (unklerhaukus):

can you make 127 simpler?

OpenStudy (unklerhaukus):

i can not see the trick to this yet

OpenStudy (anonymous):

@heena plzz help

OpenStudy (unklerhaukus):

well it turns out that 127 is a prime number

OpenStudy (unklerhaukus):

i dont know if this helps or not

OpenStudy (anonymous):

@heena can u solve this......

OpenStudy (anonymous):

wait lemme see i didnt read qn yet lol :P

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

it so long :O @Ishaan94 u have scientific calculator plz do dis..

OpenStudy (anonymous):

i think there is a trick for doing this

OpenStudy (unklerhaukus):

the product has over 170 places, a calculator will not help here

OpenStudy (anonymous):

7^1 = 7 7^2 = 49 7^3 = ...3 7^4 = ...1 7^5 = ...7 2^1 = 2 2^2 = 4 2^3 = 8 2^4 = 16 2^5 = 32 127^79 = ...1? 2^24= ...6? So, 6?

OpenStudy (anonymous):

The answer is 8 plzz check it

OpenStudy (anonymous):

Ohh hmm I must have done silly mistake

OpenStudy (anonymous):

Yes, 127^79= ...3 6x3 = 18 and 8 is the answer

OpenStudy (anonymous):

can u tell the tricks used..

OpenStudy (unklerhaukus):

i do not understand how you did that @Ishaan94

OpenStudy (anonymous):

Hmm 23*24 = ...2, we only need to get multiplication of the digits in unit place. so, to get unit place digit for 127^(79)*2^(24), we need get unit place of 7^(79)*2^24

OpenStudy (unklerhaukus):

yeah i get that bit , but i don not get how you found the unit place of 7^{79}=....3

OpenStudy (anonymous):

how u get the unit digit for 2^24 as 6 and other as 3//

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

their unit digits repeat after certain number of multiplication. 7^1 = 7 7^2 = 49 7^3 = 343 7^4 = 2401 7^5 = 16807 7^6 = 117649 7^7 = 823453 7^8 = 5 764 801 try to see the pattern you will realize. i am not a good teacher, sorry.

OpenStudy (anonymous):

so.....

OpenStudy (unklerhaukus):

i think i see what you are getting at @Ishaan94

OpenStudy (anonymous):

after 4 powers, 7 gets back to 1st form. 79 = 80 -1 => 79=4*20 - 1. so, 7^79 must be of ...3 form

OpenStudy (unklerhaukus):

i didn't know there were such patterns for powers

OpenStudy (anonymous):

then for 2^24

OpenStudy (anonymous):

Try it yourself. :-) 4,8,16,32,...

OpenStudy (anonymous):

so it repeates after 6 so 2^24 must be 6 is that the way

OpenStudy (anonymous):

No not 6 but 4 2,4,8,16,32,64,128,256,512,...

OpenStudy (anonymous):

hey it repeates after 16

OpenStudy (anonymous):

it does, but the actual interval is 4 not 16.

OpenStudy (anonymous):

ok

OpenStudy (unklerhaukus):

1,3,9,27,81,243,279,2197,..1,..3,..9,..7,..1,..3,...9,...7,......

OpenStudy (unklerhaukus):

kinda reminds me of repeating decimals

OpenStudy (anonymous):

I got it))) http://www.wolframalpha.com/input/?i=127%5E79+*++2%5E24 The last digit in the resultant number is 8.